OpenStudy (anonymous):

Simplify. 14 – (–8) + 5 – (–10) A. –9 B. 1 C. 21 D. 37

OpenStudy (anonymous):

@jdoe0001 @whpalmer4

OpenStudy (anonymous):

Just remember one thing for this one and it will be very easy. Any time you see -(-) Scratch out both -'s and write a +

OpenStudy (anonymous):

can you give me the answer because it's due in 10 minutes and I have like 7 more questions I'll fan and medal

OpenStudy (anonymous):

No. Please read the Code of Conduct. (link at bottom of screen) We are here to help you learn interactively how to solve problems on your own. Do you understand what I said about changing all -- 's to + ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

In your problem...do you you see any double minuses?

OpenStudy (anonymous):

@Yungwisdom2000 can you help me because I don't get this I'm sorry to keep bugging you @whpalmer4

OpenStudy (anonymous):

um was I supposed too?

OpenStudy (anonymous):

Mods...Close this thread please.

OpenStudy (anonymous):

whats up

OpenStudy (anonymous):

I don't get this at all... I need help

OpenStudy (whpalmer4):

\[14 - (-8) =\] Think of subtraction of a positive number as moving to the left on the number line. Subtract, move 1 unit to the left. Subtract 2, move 2 units to the left. If you agree with that, what would subtracting -1 do?

OpenStudy (anonymous):

@whpalmer4.... This person doesn't want to know how to do it. Only wants answers. Did you not read above?

OpenStudy (anonymous):

the answer is 22 use a calculator

OpenStudy (anonymous):

I don't mind the help though

OpenStudy (whpalmer4):

Using a calculator is the wrong thing to do if you don't understand what is going on! Using a calculator just results in more mistakes that you don't notice because you don't have a feel for the underlying numbers, and people who need calculators tend (unfortunately) to not be very good at using them correctly!

OpenStudy (whpalmer4):

So, @ShaTay if moving 1 spot to the left on the number line corresponds to subtracting 1, what would you do if subtracting -1?

OpenStudy (anonymous):

I got my answer... it's 37 right?

OpenStudy (anonymous):

I didn't

OpenStudy (whpalmer4):

yes. when you see something like \[x- (-y)\]you can replace that with \[x +y\] so the equation becomes \[14 - (-8) + 5 - (-10) = 14 + 8 + 5 + 10 = 22 + 15 = 37\]

OpenStudy (anonymous):

right thank you